Write a recursive formula for the following sequence: 7, 10, 21, 30, 63, 90, 189, 270
Answer by richwmiller(17219)

7, 10, 21, 30, 63, 90, 189, 270, 567, 810, 1701, 2430, 5103, 7290, 15309, 21870, 45927, 65610, 137781, 196830, 413343, 590490, 1240029, 1771470, 3720087, 5314410, 11160261, 15943230, 33480783, 47829690, 100442349, 143489070, 301327047, 430467210, 903981141, 1291401630, 2711943423, 3874204890, 8135830269, 11622614670, 24407490807, 34867844010, 73222472421, 104603532030, 219667417263, 313810596090, 659002251789, 941431788270, 1977006755367, 2824295364810, 5931020266101, 8472886094430, 17793060798303, 25418658283290, 53379182394909, 76255974849870, 160137547184727, 228767924549610, 480412641554181, 686303773648830, ...

7+3^1=10
3^1*7=21
3^2+7*3^1=30
7*3^2=63
or
3^3+3^1*7=90
7*3^3=189
(-7-10z)/(3z^2-1)
or
An = 1/2 3^(n/2-3/2)(21-21(-1)^n+10sqrt(3)+10 (-1)^n sqrt(3))